OpenStudy (anonymous):

solve my substitution (medal and fan) -x - y - z = -8 -4x + 4y + 5z = 7 2x + 2z = 4

OpenStudy (texaschic101):

2x + 2z = 4 2x = -2z + 4 x = -z + 2 -x - y - z = -8 -(-z + 2) - y - z = -8 z - 2 - y - z = -8 -2 - y = -8 -y = -8 + 2 -y = -6 y = 6 -4x + 4y + 5z = 7 -4(-z + 2) + 4(6) + 5z = 7 4z - 8 + 24 + 5z = 7 9z + 16 = 7 9z = 7 - 16 9z = - 9 z = -1 x = -z + 2 x = -(-1) + 2 x = 1 + 2 x = 3 lets check it now... -x - y - z = -8 -3 - 6 -(-1) = -8 -9 + 1 = -8 -8 = -8 (correct) check.. -4x + 4y + 5z = 7 -4(3) + 4(6) + 5(-1) = 7 -12 + 24 - 5 = 7 - 17 + 24 = 7 7 = 7 (correct) your answer is : x = 3, y = 6, and z = -1 or (3,6,-1)